1 -2 Find the general solution to the homogeneous system of DE: 3 2 6 x = Ax where A = -2 1 -2
Answer #1

A 3 2 2 6 - -2 -2 - 4 the characteristic equation in 3-7 -2 2 6 1-7 2 -u-y = 0 - 2 -1 »)(2+3)=4-4) -2(2x+8-2)+b(4+1-4)= 0 3 -